# On Creating a Causally Grounded Usable Rating Method for Assessing the Robustness of Foundation Models Supporting Time Series

### v2

@@ Appendix G Source Code for Data Processing
53 plt.close()
Appendix H Additional Implementation Details
All Forecasting Model Training Systems (FMTS) were executed on Colab notebooks utilizing the L4 GPU available through Colab Pro, which offers 22.5 GB of GPU RAM. Additional details regarding the models such as the inference times and other architectural details can be found in Section 5.1.
Hyperparameters set
  • MOMENT: head_dropout: 0.1, weight_decay: 0, freeze_encoder: True, freeze_embedder: True, freeze_head: False
  • Phi-3: _attn_implementation=’eager’, max_new_tokens: 300, temperature: 0.0, do_sample: False
  • Gemini: Temperature: 0. Rest of the parameters were default.
Appendix I Reproducibility Checklist
This paper:
  1. 1.
    Includes a conceptual outline and/or pseudocode description of AI methods introduced
    Answer: Yes (Appendix A).
  2. 2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results
    Answer: Yes (Section 5.2)
  3. 3.
    Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper
    Answer: Yes (Sections 1, 2)
    Does this paper make theoretical contributions? (yes/no)
    Answer: No
  4. 4.
    Does this paper rely on one or more datasets?
    Answer: Yes (Description in Section 5.1
  5. 5.
    A motivation is given for why the experiments are conducted on the selected datasets
    Answer: Yes (Sections 1, 2)
  6. 6.
    All novel datasets introduced in this paper are included in a data appendix.
    Answer: NA (We used an existing dataset from Yahoo! Finance.)
  7. 7.
    All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes.
    Answer: NA
  8. 8.
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations.
    Answer: NA
  9. 9.
    All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available.
    Answer: Yes (Section 5.1)
  10. 10.
    All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying.
    Answer: NA
  11. 11.
    Does this paper include computational experiments?
    Answer: Yes
  12. 12.
    Any code required for pre-processing data is included in the appendix.
    Answer: Yes, code required to convert the data downloaded from Yahoo! finance to sliding window, apply perturbations, and generate time series plots are provided in Appendix G.
  13. 13.
    All source code required for conducting and analyzing the experiments is included in a code appendix.
    Answer: Yes. You can find the source code and data here: https://anonymous.4open.science/r/rating-fmts-1B30/README.md
  14. 14.
    All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes.
    Answer: Yes
  15. 15.
    All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from
    Answer: Yes
  16. 16.
    If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results.
    Answer: Yes (provided in the source code).
  17. 17.
    This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks.
    Answer: Yes (in Appendix H)
  18. 18.
    This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics.
    Answer: Yes (Section 4.1 and Appendix C)
  19. 19.
    This paper states the number of algorithm runs used to compute each reported result.
    Answer: No.
  20. 20.
    Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information.
    Answer: Yes (Section 5.2, Appendix E, Section 5.4, and Appendix F)
  21. 21.
    The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank).
    Answer: Yes (Section 5.2, Appendix E, Section 5.4, and Appendix F)
  22. 22.
    This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments. (yes/partial/no/NA)
    Answer: Yes (Appendix Appendix H).
  23. 23.
    This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting. (yes/partial/no/NA)
    Answer: NA

### v3

@@ 5.4. User Study
To validate HP3, we used the Spearman Rank Correlation coefficient (Zar, 2005) ($\rho$) to evaluate the alignment between the users’ rankings and those produced by our approach. We considered a confidence interval of 90 %. The fairness panel showed a high correlation ($\rho$= 0.73), and the robustness under P1 showed a strong correlation ($\rho$ = 0.91). However, robustness under P2 showed a weak correlation ($\rho$ = 0.14).
In summary, the results of the user study indicate that the rankings generated by our approach can significantly reduce the difficulty of comparing the robustness of different TSFM systems. However, when the comparison metric is fairness, this reduction is not significant. Additionally, while user rankings align well with our method generated rankings for fairness and robustness under P1, they show a weak correlation for robustness under P2, indicating differing perceptions of the ’value halved’ perturbation. P1 (drop-to-zero) involves a significant semantic change that is easier to spot, whereas P2 (value halved) is also a semantic perturbation but subtler, making it potentially harder to identify. Our current study is preliminary and promising; an avenue for future work is to conduct it at a larger scale.
Forecasting Evaluation Dimensions P Partial Order Complete Order
[…]
P1 $S_{a}$: 2.6, $S_{m}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{r}$: 4.6, $S_{p}^{ni}$: 5.9, $S_{v1}$: 5.9, $S_{p}$: 6.9, $S_{c}$: 6.9, $S_{v2}$: 6.9, $S_{b}$: 6.9 } $S_{a}$: 1, $S_{m}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 1, $S_{p}^{ni}$: 2, $S_{v1}$: 2, $S_{p}$: 3, $S_{c}$: 3, $S_{v2}$: 3, $S_{v2}$: 3, $S_{b}$: 3}
P2 $S_{a}$: 4.6, $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{m}$: 4.6, $S_{r}$: 4.6, $S_{v1}$: 5.9, $S_{c}$: 6.9, $S_{p}$: 6.9, $S_{v2}$: 6.9, $S_{b}$: 6.9} $S_{a}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{m}$: 1, $S_{r}$: 1, $S_{v1}$: 2, $S_{c}$: 3, $S_{p}$: 3, $S_{v2}$: 3, $S_{b}$: 3}
P4 $S_{v2}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{r}$: 4.6, $S_{v1}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{b}$: 6.9 } $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{b}$: 3 }
P5 $S_{v2}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{r}$: 5.2, $S_{v1}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{b}$: 6.9} $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{b}$: 3}
P6 $S_{r}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{v2}$: 5.2, $S_{v1}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{b}$: 6.9} $S_{r}$: 1, $S_{g}^{ni}$: 1, $S_{v2}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{b}$: 2}
[…]
P1 $S_{a}$: 0.6, $S_{c}$: 4.6, $S_{v1}$: 4.6, $S_{v2}$: 4.6, $S_{p}$: 5.9, $S_{p}^{ni}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{r}$: 5.9, $S_{g}$: 6.9, $S_{m}$: 6.9, $S_{b}$: 6.9 } $S_{a}$: 1, $S_{c}$: 1, $S_{v1}$: 1, $S_{v2}$: 1, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{g}^{ni}$: 2, $S_{r}$: 2, $S_{g}$: 3, $S_{m}$: 3, $S_{b}$: 3 }
P2 $S_{a}$: 2.6, $S_{c}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{v1}$: 4.6, $S_{r}$: 4.6, $S_{v2}$: 5.2, $S_{p}$: 5.2, $S_{g}$: 5.9, $S_{g}^{ni}$: 5.9, $S_{m}$: 6.9, $S_{b}$: 6.9 } $S_{a}$: 1, $S_{c}$: 1, $S_{p}^{ni}$: 1, $S_{v1}$: 1, $S_{r}$: 1, $S_{v2}$: 2, $S_{p}$: 2, $S_{g}$: 2, $S_{g}^{ni}$: 2, $S_{m}$: 3, $S_{b}$: 3 }
P3 $S_{g}$: 4.6, $S_{g}^{ni}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{p}$: 4.6, $S_{c}$: 4.6, $S_{m}$: 6.9, $S_{a}$: 6.9, $S_{r}$: 6.9, $S_{b}$: 6.9} $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{p}$: 1, $S_{c}$: 1, $S_{m}$: 2, $S_{a}$: 2, $S_{r}$: 2, $S_{b}$: 2}
P4 $S_{v2}$: 4.6, $S_{v1}$: 5.9, $S_{p}^{ni}$: 6.9, $S_{g}^{ni}$: 6.9, $S_{r}$: 6.9, $S_{b}$: 6.9 } $S_{v2}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 3, $S_{g}^{ni}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P5 $S_{v2}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{v1}$: 5.2, $S_{r}$: 5.9, $S_{g}^{ni}$: 6.9, $S_{b}$: 6.9 } $S_{v2}$: 1, $S_{p}^{ni}$: 1, $S_{v1}$: 1, $S_{r}$: 2, $S_{g}^{ni}$: 3, $S_{b}$: 3 }
P6 $S_{r}$: 4.6, $S_{p}^{ni}$: 4.6, $S_{v2}$: 5.9, $S_{v1}$: 6.9, $S_{g}^{ni}$: 6.9, $S_{b}$: 6.9 } $S_{r}$: 1, $S_{p}^{ni}$: 1, $S_{v2}$: 2, $S_{v1}$: 3, $S_{g}^{ni}$: 3, $S_{b}$: 3 }
Confounding Bias with Industry as confounder (PIEI $\downarrow$) P1 $S_{g}^{ni}$: 437.28, $S_{v1}$: 539.98, $S_{v2}$: 609.44, $S_{g}$: 1012.79, $S_{a}$: 1223.59, $S_{p}$: 1586.51, $S_{p}^{ni}$: 2003.93, $S_{r}$: 4109.55, $S_{c}$: 4111.33, $S_{m}$: 4176.43, $S_{b}$: 5237.52 } $S_{g}^{ni}$: 1, $S_{v1}$: 1, $S_{v2}$: 1, $S_{g}$: 1, $S_{a}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 2, $S_{c}$: 3, $S_{m}$: 3, $S_{b}$: 3 }
P2 $S_{v2}$: 539.12, $S_{a}$: 1531.37, $S_{g}^{ni}$: 2086.5, $S_{p}^{ni}$: 2211.66, $S_{m}$: 2536.07, $S_{p}$: 3071.64, $S_{g}$: 3140.33, $S_{c}$: 3909.98, $S_{v1}$: 4405.36, $S_{r}$: 4681.26, $S_{b}$: 9593.46} $S_{v2}$: 1, $S_{a}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{m}$: 2, $S_{p}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{v1}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P3 $S_{g}$: 349.16, $S_{g}^{ni}$: 493.42, $S_{c}$: 1036.66, $S_{p}^{ni}$: 1247.39, $S_{a}$: 2251.9, $S_{p}$: 3043.22, $S_{r}$: 4509.75, $S_{b}$: 5225.69, $S_{m}$: 5715.21} $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{c}$: 1, $S_{p}^{ni}$: 2, $S_{a}$: 2, $S_{p}$: 2, $S_{r}$: 3, $S_{b}$: 3, $S_{m}$: 3}
P4 $S_{v1}$: 450.09, $S_{p}^{ni}$: 1372.59, $S_{g}^{ni}$: 1498.45, $S_{v2}$: 1941.72, $S_{r}$: 3916.46, $S_{b}$: 6746.66} $S_{v1}$: 1, $S_{p}^{ni}$: 1, $S_{g}^{ni}$: 2, $S_{v2}$: 2, $S_{r}$: 3, $S_{b}$: 3}
P5 $S_{g}^{ni}$: 2391.58, $S_{v1}$: 2571.56, $S_{v2}$: 4128.95, $S_{r}$: 5004.2, $S_{b}$: 7883.56, $S_{p}^{ni}$: 13472.24} $S_{g}^{ni}$: 1, $S_{v1}$: 1, $S_{v2}$: 2, $S_{r}$: 2, $S_{b}$: 3, $S_{p}^{ni}$: 3}
P6 $S_{v2}$: 205.36, $S_{g}^{ni}$: 789.78, $S_{p}^{ni}$: 1687.56, $S_{r}$: 2586.99, $S_{v1}$: 4215.45, $S_{b}$: 9606.06} $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 2, $S_{r}$: 2, $S_{v1}$: 3, $S_{b}$: 3}
Confounding Bias with Company as confounder (PIEC $\downarrow$) P1 $S_{g}^{ni}$: 632.31, $S_{a}$: 647.15, $S_{v2}$: 761.01, $S_{g}$: 882.66, $S_{m}$: 910.47, $S_{p}$: 914.62, $S_{p}^{ni}$: 993.88, $S_{r}$: 1460.55, $S_{c}$: 1630.63, $S_{b}$: 1855.99, $S_{v1}$: 1923.8} $S_{g}^{ni}$: 1, $S_{a}$: 1, $S_{v2}$: 1, $S_{g}$: 1 $S_{m}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 2, $S_{c}$: 3, $S_{b}$: 3, $S_{v1}$: 3 }
P2 $S_{r}$: 295.4, $S_{v1}$: 516.73, $S_{g}^{ni}$: 709.9, $S_{g}$: 829.27, $S_{v2}$: 984.88, $S_{m}$: 1119.67, $S_{a}$: 1234.68, $S_{b}$: 2132.45, $S_{p}$: 2566.84, $S_{p}^{ni}$: 4767.08, $S_{c}$: 4963.77} $S_{r}$: 1, $S_{v1}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{v2}$: 2, $S_{m}$: 2, $S_{a}$: 2, $S_{b}$: 2, $S_{p}$: 3, $S_{p}^{ni}$: 3, $S_{c}$: 3}
P3 $S_{p}^{ni}$: 429.43, $S_{g}$: 476.25, $S_{m}$: 807.34, $S_{c}$: 880.64, $S_{g}^{ni}$: 938.41, $S_{a}$: 1801.04, $S_{r}$: 2051.79, $S_{b}$: 2201.23, $S_{p}$: 3572.98} $S_{p}^{ni}$: 1, $S_{g}$: 1, $S_{m}$: 1, $S_{c}$: 2, $S_{g}^{ni}$: 2, $S_{a}$: 2, $S_{r}$: 3, $S_{b}$: 3, $S_{p}$: 3}
P4 $S_{g}^{ni}$: 363.8, $S_{p}^{ni}$: 760.76, $S_{r}$: 1038.87, $S_{v2}$: 1584.25, $S_{v1}$: 2122.6, $S_{b}$: 2635.14} $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{v2}$: 2, $S_{v1}$: 3, $S_{b}$: 3}
P5 $S_{g}^{ni}$: 960.09, $S_{v1}$: 1058.15, $S_{r}$: 1211.5, $S_{p}^{ni}$: 1522.82, $S_{v2}$: 1596.87, $S_{b}$: 1952.83} $S_{g}^{ni}$: 1, $S_{v1}$: 1, $S_{r}$: 2, $S_{p}^{ni}$: 2, $S_{v2}$: 3, $S_{b}$: 3}
P6 $S_{v1}$: 678.24, $S_{v2}$: 1112.99, $S_{r}$: 1155.87, $S_{g}^{ni}$: 1990.4, $S_{b}$: 2027.3, $S_{p}^{ni}$: 2867.36} $S_{v1}$: 1, $S_{v2}$: 1, $S_{r}$: 2, $S_{g}^{ni}$: 2, $S_{b}$: 3, $S_{p}^{ni}$: 3}
[…]
P2 $S_{g}$: 2.69, $S_{v1}$: 3.77, $S_{p}^{ni}$: 4.95, $S_{v2}$: 6.04, $S_{g}^{ni}$: 9.12, $S_{c}$: 9.54, $S_{p}$: 13.09, $S_{a}$: 16.38, $S_{m}$: 16.58, $S_{r}$: 25.52, $S_{b}$: 101.14 } $S_{g}$: 1, $S_{v1}$: 1, $S_{p}^{ni}$: 1, $S_{v2}$: 1, $S_{g}^{ni}$: 2, $S_{c}$: 2, $S_{p}$: 2, $S_{a}$: 2, $S_{m}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P3 $S_{m}$: 3.6, $S_{g}$: 4.88, $S_{g}^{ni}$: 5.31, $S_{a}$: 7.5, $S_{c}$: 11.51, $S_{p}$: 14.8, $S_{r}$: 15.0, $S_{p}^{ni}$: 16.25, $S_{b}$: 51.42} $S_{m}$: 1, $S_{g}$: 1, $S_{g}^{ni}$: 1, $S_{a}$: 2, $S_{c}$: 2, $S_{p}$: 2, $S_{r}$: 3, $S_{p}^{ni}$: 3, $S_{b}$: 3}
P4 $S_{v2}$: 4.67, $S_{g}^{ni}$: 5.21, $S_{v1}$: 6.38, $S_{p}^{ni}$: 16.06, $S_{r}$: 41.37, $S_{b}$: 53.09} $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3}
P5 $S_{g}^{ni}$: 4.6, $S_{v2}$: 5.24, $S_{v1}$: 7.9, $S_{p}^{ni}$: 10.74, $S_{r}$: 55.88, $S_{b}$: 98.69} $S_{g}^{ni}$: 1, $S_{v2}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3}
P6 $S_{v2}$: 1.06, $S_{g}^{ni}$: 1.15, $S_{v1}$: 9.15, $S_{p}^{ni}$: 18.32, $S_{r}$: 27.08, $S_{b}$: 52.01} $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{v1}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3}
[…]
P2 $S_{b}$: 0.0, $S_{p}^{ni}$: 1.94, $S_{c}$: 3.94, $S_{v2}$: 3.94, $S_{g}$: 4.33, $S_{v1}$: 6.14, $S_{r}$: 6.17, $S_{m}$: 7.78, $S_{g}^{ni}$: 8.14, $S_{a}$: 13.25, $S_{p}$: 19.38 } $S_{b}$: 1, $S_{p}^{ni}$: 1, $S_{c}$: 1, $S_{v2}$: 1, $S_{g}$: 2, $S_{v1}$: 2, $S_{r}$: 2, $S_{m}$: 2, $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{p}$: 3 }
P3 $S_{b}$: 0.0, $S_{a}$: 2.5, $S_{c}$: 4.78, $S_{g}$: 5.48, $S_{r}$: 6.13, $S_{p}$: 6.17, $S_{p}^{ni}$: 6.25, $S_{m}$: 8.36, $S_{g}^{ni}$: 9.3} $S_{b}$: 1, $S_{a}$: 1, $S_{c}$: 1, $S_{g}$: 2, $S_{r}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 3, $S_{m}$: 3, $S_{g}^{ni}$: 3}
P4 $S_{b}$: 0.0, $S_{v2}$: 2.11, $S_{v1}$: 3.36, $S_{g}^{ni}$: 5.5, $S_{r}$: 9.49, $S_{p}^{ni}$: 11.09} $S_{b}$: 1, $S_{v2}$: 1, $S_{v1}$: 2, $S_{g}^{ni}$: 2, $S_{r}$: 3 $S_{p}^{ni}$: 3}
P5 $S_{b}$: 0.0, $S_{v2}$: 3.96, $S_{v1}$: 5.52, $S_{g}^{ni}$: 6.67, $S_{p}^{ni}$: 7.65, $S_{r}$: 13.53} $S_{b}$: 1, $S_{v2}$: 1, $S_{v1}$: 2, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 3, $S_{r}$: 3}
P6 $S_{b}$: 0.0, $S_{v2}$: 1.03, $S_{g}^{ni}$: 4.18, $S_{r}$: 4.49, $S_{p}^{ni}$: 7.59, $S_{v1}$: 20.31} $S_{b}$: 1, $S_{v2}$: 1, $S_{g}^{ni}$: 2, $S_{r}$: 2, $S_{p}^{ni}$: 3, $S_{v1}$: 3}
Table 5. Final raw scores and ratings based on different metrics computed. Higher rating indicate higher bias.
Forecasting Evaluation Dimensions P Partial Order Complete Order
[…]
P1 $S_{v1}$: 0.064, $S_{c}$: 0.065, $S_{g}^{ni}$: 0.067, $S_{g}$: 0.072, $S_{a}$: 0.084, $S_{m}$: 0.100, $S_{p}$: 0.100, $S_{p}^{ni}$: 0.100, $S_{v2}$: 0.127, $S_{r}$: 0.830, $S_{b}$: 1.276 } $S_{v1}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 2, $S_{a}$: 2, $S_{m}$: 2, $S_{p}$: 2, $S_{p}^{ni}$: 2, $S_{v2}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P2 $S_{v1}$: 0.047, $S_{g}$: 0.051, $S_{c}$: 0.053, $S_{g}^{ni}$: 0.060, $S_{v2}$: 0.068, $S_{a}$: 0.069, $S_{p}^{ni}$: 0.095, $S_{m}$: 0.098, $S_{p}$: 0.100, $S_{r}$: 0.830, $S_{b}$: 1.276 } $S_{v1}$: 1, $S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{v2}$: 2, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3 }
P4 $S_{v1}$: 0.039, $S_{v2}$: 0.041, $S_{g}^{ni}$: 0.047, $S_{p}^{ni}$: 0.080, $S_{r}$: 0.830, $S_{b}$: 1.276 } $S_{v1}$: 1, $S_{v2}$: 1, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P6 $S_{v2}$: 0.041, $S_{g}^{ni}$: 0.047, $S_{p}^{ni}$: 0.079, $S_{v1}$: 0.089, $S_{r}$: 0.832, $S_{b}$: 1.276 } $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 2, $S_{v1}$: 2, $S_{r}$: 3, $S_{b}$: 3 }
[…]
P1 $S_{v1}$: 5.30, $S_{c}$: 5.40, $S_{g}^{ni}$: 5.65, $S_{g}$: 6.13, $S_{p}^{ni}$: 8.87, $S_{p}$: 9.19, $S_{m}$: 9.32, $S_{v2}$: 11.18, $S_{a}$: 18.36, $S_{r}$: 86.99, $S_{b}$: 947.56} $S_{v1}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{g}$: 1, $S_{p}^{ni}$: 2, $S_{p}$: 2, $S_{m}$: 2, $S_{v2}$: 2, $S_{a}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P2 $S_{v1}$: 4.24, $S_{g}$: 4.74, $S_{c}$: 4.99, $S_{g}^{ni}$: 5.59, $S_{v2}$: 6.16, $S_{a}$: 8.24, $S_{p}^{ni}$: 8.49, $S_{m}$: 9.15, $S_{p}$: 9.32, $S_{r}$: 86.87, $S_{b}$: 947.56} $S_{v1}$: 1, $S_{g}$: 1, $S_{c}$: 1, $S_{g}^{ni}$: 1, $S_{v2}$: 2, $S_{a}$: 2, $S_{p}^{ni}$: 2, $S_{m}$: 2, $S_{p}$: 3, $S_{r}$: 3, $S_{b}$: 3}
P4 $S_{v1}$: 3.68, $S_{v2}$: 3.89, $S_{g}^{ni}$: 4.46, $S_{p}^{ni}$: 7.10, $S_{r}$: 86.65, $S_{b}$: 947.56} $S_{v1}$: 1, $S_{v2}$: 1, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3}
P5 $S_{v1}$: 3.67, $S_{v2}$: 3.90, $S_{g}^{ni}$: 4.47, $S_{p}^{ni}$: 7.23, $S_{r}$: 86.53, $S_{b}$: 947.56} $S_{v1}$: 1, $S_{v2}$: 1, $S_{g}^{ni}$: 2, $S_{p}^{ni}$: 2, $S_{r}$: 3, $S_{b}$: 3}
P6 $S_{v2}$: 3.93, $S_{g}^{ni}$: 4.42, $S_{p}^{ni}$: 7.24, $S_{v1}$: 8.26, $S_{r}$: 87.20, $S_{b}$: 947.56} $S_{v2}$: 1, $S_{g}^{ni}$: 1, $S_{p}^{ni}$: 2, $S_{v1}$: 2, $S_{r}$: 3, $S_{b}$: 3}
Accuracy (Sign Accuracy $\uparrow$) P0 $S_{m}$: 40.70, $S_{p}$: 45.09, $S_{p}^{ni}$: 47.67, $S_{r}$: 49.88, $S_{g}^{ni}$: 50.41, $S_{v2}$: 51.28, $S_{v1}$: 51.32, $S_{g}$: 52.08, $S_{c}$: 53.75, $S_{a}$: 60.08, $S_{b}$: 62.60 } $S_{m}$: 1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 1, $S_{g}^{ni}$: 2, $S_{v2}$: 2, $S_{v1}$: 2, $S_{g}$: 2, $S_{c}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
P1 $S_{m}$: 41.19, $S_{v2}$: 41.54, $S_{p}$: 44.33, $S_{p}^{ni}$: 46.77, $S_{v1}$: 48.77, $S_{r}$: 49.62, $S_{g}$: 50.53, $S_{c}$: 52.09, $S_{g}^{ni}$: 53.93, $S_{a}$: 57.08, $S_{b}$: 62.60 } $S_{m}$: 1, $S_{v2}$:1, $S_{p}$: 1, $S_{p}^{ni}$: 1, $S_{v1}$: 2, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2 $S_{g}^{ni}$: 3, $S_{a}$: 3, $S_{b}$: 3 }
P2 $S_{m}$: 41.05, $S_{p}$: 44.02, $S_{v2}$: 45.28, $S_{p}^{ni}$: 47.67, $S_{r}$: 49.64, $S_{g}$: 49.75, $S_{c}$: 50.79, $S_{g}^{ni}$: 54.43, $S_{a}$: 57.13, $S_{v1}$: 58.69, $S_{b}$: 62.60 } $S_{m}$: 1, $S_{p}$: 1, $S_{v2}$: 1, $S_{p}^{ni}$: 1, $S_{r}$: 2, $S_{g}$: 2, $S_{c}$: 2, $S_{g}^{ni}$: 2, $S_{a}$: 3, $S_{v1}$: 3, $S_{b}$: 3 }
P4 $S_{p}^{ni}$: 42.87, $S_{r}$: 49.71, $S_{g}^{ni}$: 49.46, $S_{v1}$: 51.35, $S_{v2}$: 54.74, $S_{b}$: 62.60 } $S_{p}^{ni}$: 1, $S_{r}$: 1, $S_{g}^{ni}$: 2, $S_{v1}$: 2, $S_{v2}$: 3, $S_{b}$: 3 }
P5 $S_{p}^{ni}$: 41.60, $S_{g}^{ni}$: 49.63, $S_{r}$: 49.67, $S_{v2}$: 51.14, $S_{v1}$: 53.95, $S_{b}$: 62.60 } $S_{p}^{ni}$: 1, $S_{g}^{ni}$: 1, $S_{r}$: 2, $S_{v2}$: 2, $S_{v1}$: 3, $S_{b}$: 3 }
P6 $S_{p}^{ni}$: 42.60, $S_{g}^{ni}$: 48.78, $S_{v1}$: 43.97, $S_{r}$: 50.05, $S_{v2}$: 52, $S_{b}$: 62.60 } $S_{p}^{ni}$: 1, $S_{g}^{ni}$: 1, $S_{v1}$: 2, $S_{r}$: 2, $S_{v2}$: 3, $S_{b}$: 3}
Table 6. Final raw scores and ratings based on different metrics computed. Higher rating indicate higher inaccuracy. For simplicity, we denoted the raw scores for accuracy metrics using just the mean value, but standard deviation was also considered for rating. The chosen rating level, L = 3.
6. Discussion and Conclusion
Our paper aimed to measure the impact of perturbations and confounders on the outcome of TSFM using stock prices across leading companies and industries. We studied Industry and Company as confounders, motivated by the intuition that stakeholders rely on learning-based systems for stock purchase decisions and would be interested in knowing if model errors depend on stock price ranges. For example, does a model commit more errors predicting META’s stock prices compared to MRK’s? To minimize volatility effects, we performed both intra-industry and inter-industry analyses. In future, we plan to study confounders like seasonal trends and financial news. As demonstrated, we believe metrics should be selected based on the questions one wants to answer, rather than relying solely on statistical accuracy. The hypothesis testing approach from (Lakkaraju et al., 2024), adapted for our work, helped quantify biases and perturbation impacts on test systems. The perturbations used in our analysis have real-world impacts, applicable to both numeric and multi-modal data. While methods like differential evaluation can find the most impactful perturbation variations, we focused on assessing whether simple, subtle perturbations affect TSFM.
Conclusion We proposed a causally grounded empirical framework to study TSFM robustness against three input perturbations, evaluating seven state-of-the-art TSFM across six prominent stocks in three industries. Our framework’s ratings accurately assessed TSFM robustness and provided actionable insights for model selection and deployment. Experiments showed multi-modal TSFM exhibited greater robustness, while uni-modal TSFM had higher forecasting accuracy. TSFM trained on time series tasks showed better robustness and accuracy compared to general-purpose TSFM. A user study confirmed our ratings effectively convey TSFM robustness to end-users, demonstrating the framework’s real-world applicability.